Which of the following will have the highest value after 20 years, assuming the same annual interest rate of 10%?

A.

$170,000 deposited after 5 years

B.

$110,000 deposited now

C.

$20,000 deposits per year for 8 years, starting from now

D.

10 yearly deposits from now that grows at 12% per year, with $10,000 as the first deposit (i.e., 10,000 now, 11,200 in a year, 12,544 in 2 years, etc.)
Answer #1
C.

$20,000 deposits per year for 8 years, starting from now has the highest value after 20 years

Working:

A. Value after 20 years = 170000*(1+0.10)^15 = $ 7,10,132.19
B. Value after 20 years = 110000*(1+0.10)^20 = $ 7,40,024.99
C.
# 1
Value after 7 years from now = Annuity*Future Value of annuity of 1
= $     20,000.00 * 12.57947691
= $ 2,51,589.54
Future value of annuity of 1 = ((((1+i)^n)-1)/i)*(1+i)
= ((((1+0.10)^8)-1)/0.10)*(1+0.10)
= 12.57948
# 2
Value after 20 years = $ 2,51,589.54 *(1+0.10)^13
= $ 8,68,555.32
D.
Deposit number Year Cash flows Future Value of 1 Future Value of cash flows
a b c=1.10^(20-a) d=b*c
1 0 $ 10,000.00 6.727499949 $     67,275.00
2 1 $ 11,200.00 6.115909045 $     68,498.18
3 2 $ 12,544.00 5.559917313 $     69,743.60
4 3 $ 14,049.28 5.054470285 $     71,011.67
5 4 $ 15,735.19 4.594972986 $     72,302.79
6 5 $ 17,623.42 4.177248169 $     73,617.39
7 6 $ 19,738.23 3.797498336 $     74,955.88
8 7 $ 22,106.81 3.452271214 $     76,318.72
9 8 $ 24,759.63 3.138428377 $     77,706.33
10 9 $ 27,730.79 2.853116706 $     79,119.17
Total $ 7,30,548.73